You were playing fetch with your dog and accidentally 

threw the toy and it got stuck on the roof of your 

house. You leaned the ladder against your house at a 

30° angle of elevation. The foot of the ladder is 7 feet 

from the foot of the house. How long is your ladder? 



1
Expert's answer
2021-01-11T17:33:42-0500

Let "L=" the length of the ladder, "\alpha=" the angle of elevation, "H=" the height of the house, and "X=" the distance from the foot of the house to the foot of the ladder.



From the right triangle


"\cos\alpha=\dfrac{X}{L}"

Then


"L=\dfrac{X}{\cos\alpha}"

Given "\alpha=30\degree, X=7\ ft."

Substitute


"L=\dfrac{7}{\cos30\degree}=\dfrac{7}{\dfrac{\sqrt{3}}{2}}=\dfrac{14\sqrt{3}}{3}(ft)\approx8.083(ft)"

The ladder is 8.1 foot long.